OpenStudy (anonymous):

What is the value of the expression x + (–y) when x = 1.2 and y = –10.7? A. 11.9 B. –9.5 C. 9.5 D. –11.9

OpenStudy (anonymous):

@inowalst @omarbirjas

OpenStudy (anonymous):

let me check this one again

OpenStudy (omarbirjas):

just x-y

OpenStudy (anonymous):

b

OpenStudy (omarbirjas):

1.2-(-10.7)

OpenStudy (anonymous):

i think

OpenStudy (omarbirjas):

Or just 1.2+10.7

OpenStudy (omarbirjas):

its A

OpenStudy (anonymous):

the 10.7 is negative

OpenStudy (anonymous):

-10.7

OpenStudy (omarbirjas):

yea and ur subtracting, 2 negatives make a positive

OpenStudy (anonymous):

u are adding

OpenStudy (anonymous):

look at the question

OpenStudy (omarbirjas):

x+(-y)=x-y

OpenStudy (anonymous):

kk

OpenStudy (omarbirjas):

substitute and you get 1.2-(-10.7)

OpenStudy (omarbirjas):

which means 1.2+10.7=11.9

OpenStudy (anonymous):

i still got -9.5

OpenStudy (anonymous):

for both ways

OpenStudy (anonymous):

@omarbirjas

OpenStudy (anonymous):

there is the link

OpenStudy (anonymous):

check it

OpenStudy (anonymous):

type it i and see what u get, calculators don't lie unless there cheap

OpenStudy (omarbirjas):

http://prntscr.com/6nw4zz

OpenStudy (anonymous):

ah, i am bad a punns

OpenStudy (omarbirjas):

*they're

OpenStudy (anonymous):

kk